48a^3b^6c/12a^2b^4: 48a^3b^6c/12a^2b^4
Answer by checkley77(3848)

48a^3b^6c/12a^2b^4
[48/12=4]
[a^3/a^2=a^(3-2)=a^1=a]
[b^6/b^4=b^(6-4)=b^2]
4ab^2c ans.